Susie's age is three times that of her niece. When you add Susie's age to her niece's age, you get 44. How old is Aunt Susie? How old is her niece?
Write an equation that represents the situation. Explain any variable used.
Solve the equation from Part (a). Show your work. State your solution as a complete sentence.

Answer by ikleyn(52776)

.

N + 3N = 44


  4N   = 44


   N   = 44/4 = 11.


ANSWER.  The niece is 11 years old.   Aunt Susie is 3*11 = 33 years old.

Solved.